# Riemannian Time Warping: Multiple Sequence Alignment in Curved Spaces

### v1

@@ V Simulations and Experiments
This section compares the proposed RTW approach with baselines in various settings considering datasets in $\mathbb{R}^{1}$, $\mathbb{S}^{1}$, $\mathbb{S}^{3}$ and $\mathbb{S}^{2}_{++}$. We also apply RTW as a pre-processing step for learning movement primitives, on both hand-written data in $\mathbb{R}^{2}$ and within a real-robot scenario in $\mathrm{SE}(3)$ using an LBR iiwa.
RTW, TTW and NTW have been implemented in PyTorch, utilizing the Adam optimizer with a learning rate of $0.01$ and Autograd for gradient computations. The best model over 256 epochs was selected during the comparisons. For RTW, we use a neural network to model ${\mathbf{\phi}_{\theta}}(\hat{z})$ with four fully connected linear layers, ReLU activations and skip connections of size 1-512-512-1025-(N-1), as suggested in [11]. However, unlike NTW, we do not initialize the weights with zero, as preliminary trials showed that this significantly degrades the performance. Instead, we apply symmetry breaking [22], initializing weights with the Xavier initialization method [23].
We use a window size of ${\nu}=10$ for the sinc interpolation in RTW/TTW (10). For the Gaussian-weighted window loss in RTW (8), we choose window size ${\kappa}=5$ and step size ${\rho}=5$, which had neglectable impacts on the runtime. The penalty term (6) for RTW/NTW is weighted with ${\lambda}=100$.
The number of learnable parameters $\theta$ in NTW and RTW is $263680+1025({N}-1)+({N}-1)$, i.e., the number of weights and biases from all layers of the neural network, whereas it is only ${N}{K}$ for TTW due to the warping function
@@ V-B Inverted Time Warping in $\mathbb{S}^{1}$
footnotetext: Modified version with an ${N}$-dimensional cost matrix for ${N}>2$. footnotetext: Modified version utilizing (10)-(14) for data in $\mathbb{S}^{1}$. footnotetext: CPU used for the experiments is an AMD® Ryzen 5 3500U. footnotetext: GPU used for the experiments is an NVIDIA GeForce RTX 4080 Super.
Signal alignment is typically performed in an unsupervised setting, making it difficult to evaluate the result w.r.t. the globally optimal solution, as this is usually not available for a dataset with $N\gg 2$ signals [11]. Here, we mitigate this issue through an inverted time warping approach [^9] Warping functions that satisfy constraints (2a)-(2c) can also be used for creating an artificial temporal shift of a given signal. Then, optimal warping functions exist to reverse the tempral shift. , which we apply to data in $\mathbb{S}^{1}$. Exponential and logarithmic map are defined as
${\mathbf{Exp}^{\mathbb{S}^{D}}_{{\bm{e}}}\left({\bm{u}}\right)}={\bm{e}}\cos({\left\|{\bm{u}}\right\|})+\frac{{\bm{u}}}{{\left\|{\bm{u}}\right\|}}\sin({\left\|{\bm{u}}\right\|})\,,$ (16)
${\mathbf{Log}^{\mathbb{S}^{D}}_{{\bm{e}}}\left({\bm{x}}\right)}=\arccos({\bm{e}}^{T}{\bm{x}})\frac{{\bm{x}}-{\bm{e}}^{T}{\bm{x}}{\bm{e}}}{{\left\|{\bm{x}}-{\bm{e}}^{T}{\bm{x}}{\bm{e}}\right\|}}\,.$ (17)
@@
The literature does not offer time warping schemes for ${N}>2$ signals with data on Riemannian manifolds. Hence, we modify and extend QDTW (presented for ${N}=2$ and $\mathbb{S}^{3}$) following the concept of MMDDTW [15]. In addition, we change the implementation of the TTW and NTW algorithms to utilize (10)-(14). These modified algorithms are compared with RTW in two settings, with ${N}=4$ and ${N}=30$ signals. For transparent results, we measure the computation time on a high-grade GPU and a standard CPU. Furthermore, we evaluate three well-established metrics
@@
where the cosine distance $\text{dist}_{\text{cos}}=1-{\left\langle{\mathbf{{\hat{x}}}_{n}}[z],{\mathbf{{\hat{y}}}}[z]\right\rangle}$ is used within each ${\mathcal{D}_{dtw}}$ distance [7]. The Restoration Accuracy describes how effectively random warping functions are reversed to restore the original signal ${\mathbf{x}_{\text{original}}}$. The Barycenter Loss measures how accurately the mean ${\mathbf{{\hat{y}}}}$ of warped signals ${\mathbf{{\hat{x}}}_{n}}$ represents the data ${\mathbf{x}_{n}}$. The Alignment Quality indicates closeness of aligned signals ${\mathbf{{\hat{x}}}_{n}}$ to their mean ${\mathbf{{\hat{y}}}}$. Each experiment is repeated 100 times to compute the average for each metric.
@@
Metric QDTWfootnotemark: TTWfootnotemark: NTWfootnotemark: RTW
Restoration Accuracy 0.6565 0.6785 0.4355 0.2909
Barycenter Loss 0.8399 1.0347 0.6914 0.6932
Alignment Quality 0.0035 0.0341 0.1172 0.0087
@@
Metric QDTWfootnotemark: TTWfootnotemark: NTWfootnotemark: RTW
Restoration Accuracy 2.9523 1.4193 0.3273
Barycenter Loss 6.6766 4.6503 4.4798
Alignment Quality 0.1491 0.8514 0.1095
CPUfootnotemark: Runtime in [s] 18.8754 226.9687 103.1399
GPUfootnotemark: Runtime in [s] 9.4830 10.9298 10.2900
@@ V-B1 Evaluation for ${N}=4$
Table III shows the result. QDTW achieves the best alignment quality due to globally optimal Dynamic Programming. However, RTW is a close second, requiring about three orders of magnitude less computation time. Furthermore, the original signal ${\mathbf{x}_{\text{original}}}$ was restored best by RTW. NTW and RTW perform equally well w.r.t. the Barycenter Loss. The slightly longer runtime of RTW when compared to TTW is justified by its superior results.
@@ V-B2 Evaluation for ${N}=30$
The result is shown in Figure 1 and Table IV. QDTW is not capable of handling such a large number of signals in a reasonable time due to the excessive Dynamic Programming technique. RTW achieves the best result w.r.t. Alignment Quality. Even though TTW also achieves a good alignment, the other two metrics reveal that the signal structure is modified. NTW and RTW are again comparable in terms of the Barycenter Loss, but RTW clearly outperforms both baselines w.r.t. the Restoration Accuracy. The benefits of using the windowed sinc interpolation become evident in the TTW and RTW runtime. Furthermore, the reduced number of parameters optimized within TTW positively affects the CPU runtime. However, on the GPU, all methods are comparably fast for ${T}=100$, due to the advantages of parallel computing.
@@ V-C Multi-dimensional Riemannian manifolds
Next, we extend our experiments to higher-dimensional data by applying RTW to the Riemannian manifolds $\mathbb{S}^{3}$ and $\mathbb{S}^{2}_{++}$.
@@ V-C1 Evaluation in $\mathbb{S}^{3}$
To further validate our results from section V-B, we apply RTW to signals of unit quaternions, as QDTW was originally proposed for data in $\mathbb{S}^{3}$. Similar to [12], we artificially generate ${N}=4$ valid signals of length ${T}=200$ from the LASA [25] dataset. This involves projecting handwritten motion data from $\mathbb{R}^{2}$ onto the unit sphere of $\mathbb{S}^{3}$.
Table V compares QDTW and RTW. We observe that the mean signal obtained from RTW is an improved representation of the original data compared to QDTW. Additionally, RTW achieves an almost optimal alignment quality comparable to QDTW, which is consistent with our previous findings.
@@
Barycenter Loss 0.0388 0.0327
Alignment Quality 0.0365 0.0379
V-C2 Evaluation in $\mathbb{S}^{2}_{++}$
We temporally align signals from the Riemannian manifold of symmetric positive definite (SPD) $2\times 2$ matrices. The exponential and logarithmic maps are
${\mathbf{Exp}^{\mathbb{S}^{D}_{++}}_{{\bm{e}}}\left({\bm{u}}\right)}=\sqrt{{\bm{e}}}\exp{\left(\sqrt{{\bm{e}}}^{-1}{\bm{u}}\sqrt{{\bm{e}}}^{-1}\right)}\sqrt{{\bm{e}}},$ (19)
${\mathbf{Log}^{\mathbb{S}^{D}_{++}}_{{\bm{e}}}\left({\bm{x}}\right)}=\sqrt{{\bm{e}}}\log{\left(\sqrt{{\bm{e}}}^{-1}{\bm{x}}\sqrt{{\bm{e}}}^{-1}\right)}\sqrt{{\bm{e}}},$ (20)
where $\sqrt{{\bm{e}}}$ and $\sqrt{{\bm{e}}}^{-1}$ denote the matrix square root and its inverse, respectively. Matrices ${\bm{x}}$ and ${\bm{u}}$ have equal dimensions.
We compare RTW with two modified versions of MMDDTW. The first version, denoted as $\text{MMDDTW}_{{\mathcal{M}}}$, computes the distance between two SPD matrices according to the geodesic distance on ${\mathcal{M}}$ defined in (14) with the Frobenius norm ${\left\|\cdot\right\|}_{F}$. The second version, denoted as $\text{MMDDTW}_{\operatorname{chol}}$, utilizes the Cholesky decomposition $\operatorname{chol}\left(\cdot\right)$ to linearize the manifold, and computes a distance through
@@ V-C2 Evaluation in $\mathbb{S}^{2}_{++}$
The data for this experiment consists of manipulability ellipsoids, which are artificially generated from the LASA [25] dataset as in [26]. A planar robot with two joints is controlled to follow the handwritten character G as shown in Figure 2. At each time step, the manipulability is computed, resulting in a dataset of ${N}=3$ signals with a length of ${T}=200$ each.
We evaluate the Barycenter Loss (18b) and the Alignment Quality (18c) using both distance measures (14) and (21). Table VI presents the results, supporting our previous findings: RTW outperforms MMDDTW in terms of Barycenter Loss, achieving near-optimal alignment with remarkable speed.
@@
Metric $\text{MMDDTW}_{{\mathcal{M}}}$ $\text{MMDDTW}_{\operatorname{chol}}$ RTW
$\text{Barycenter Loss}_{{\mathcal{M}}}$ 1.0153 1.0370 0.9579
$\text{Barycenter Loss}_{\operatorname{chol}}$ 0.3449 0.3387 0.3285
$\text{Alignment Quality}_{{\mathcal{M}}}$ 1.0166 1.0463 1.0566
$\text{Alignment Quality}_{\operatorname{chol}}$ 0.3488 0.3389 0.3447
V-D Learning Movement Primitives with RTW pre-processing
Kinesthetic teaching provides an easy-to-use interface to learn robot motion skills. Demonstrated trajectories commonly vary in space and time, and are used to imitate the demonstrated behavior, for example using Dynamic Movement Primitives (DMP) [27], Gaussian Mixture Models (GMM) [6] or Gaussian Processes (GP). As a baseline, the mean signal ${\mathbf{{\hat{y}}}}$ may also serve as time-dependent motion skill encoding. Time warping has been utilized as a pre-processing technique for motion represented in Euclidean space (such as the robot’s joint-space) [5, 6]. However, to the best of our knowledge, related works on end-effector motion in $\mathrm{SE}(3)$ directly learn from raw data without any pre-processing, e.g., [21].
We assess the advantages of using RTW as a pre-processing step rather than relying on raw data, focusing on $\mathbb{R}^{2}$ in this subsection and addressing $\mathrm{SE}(3)$ afterwards. Given ${N}=4$ signals of length ${T}=200$ from the LASA dataset [25], we perform leave one signal-out cross validation.
TABLE VII: Reproduction error of unseen signals with Imitation Learning
DTW distance MAE

### v3

@@ V Simulations and Experiments
This section compares the proposed RTW approach with baselines in various settings considering datasets in $\mathbb{R}^{D}$, $\mathbb{S}^{D}$ and $\mathbb{S}^{D}_{++}$. We also apply RTW as a pre-processing step for learning movement primitives for a real-robot teaching scenario in $\mathbb{R}^{3}\times\mathbb{S}^{3}$ using an LBR iiwa manipulator.
RTW, TTW and NTW have been implemented in PyTorch, utilizing the Adam optimizer with a learning rate of $0.01$ and Autograd for gradient computations. The best model over 256 epochs was selected during the comparisons. For RTW, we use a neural network to model ${\mathbf{\phi}_{\theta}}(\hat{z})$ with four fully connected linear layers666 We also evaluated simpler model architectures, such as estimating coefficients of a polynomial, which achieved comparable results at faster runtime for our experiments. However, for generalization purposes, and to maintain a fair comparison with NTW, we stick to the neural network architecture. , ReLU activations and skip connections of size 1-512-512-1025-(N-1), as suggested in [11]. However, unlike NTW, we do not initialize the weights with zero, as preliminary trials have shown that this significantly degrades the performance. Instead, we apply symmetry breaking [24], initializing weights with the Xavier initialization method [25].
We use a window size of ${\nu}=10$ for the sinc interpolation in RTW/TTW (10). For the Gaussian-weighted window loss in RTW (8), we choose window size ${\kappa}=5$ and step size ${\rho}=5$, which had negligible impacts on the runtime. The penalty term (6) for RTW/NTW is weighted with ${\lambda}=100$.
The number of learnable parameters $\theta$ in NTW and RTW is $263680+1025({N}-1)+({N}-1)$, i.e., the number of weights and biases from all layers of the neural network, whereas it is only ${N}{K}$ for TTW due to the design of the warping function
@@ V-B Inverted Time Warping in $\mathbb{S}^{1}$
footnotetext: Modified version utilizing (14) for data on Riemannian manifolds. footnotetext: Modified version utilizing (10)-(14) for data on Riemannian manifolds. footnotetext: CPU used for the experiments is an AMD® Ryzen 5 3500U. footnotetext: GPU used for the experiments is an NVIDIA GeForce RTX 4080 Super.
Signal alignment is typically performed in an unsupervised setting, making it difficult to evaluate the result w.r.t. the optimal solution, as it is usually not available for datasets with $N\gg 2$ signals [11]. Here, we mitigate this issue through an inverted time warping approach [^11] Warping functions that satisfy constraints (2a)-(2c) can also be used for creating an artificial temporal shift of a given signal. Then, optimal warping functions exist to reverse the tempral shift. , which we apply to data in $\mathbb{S}^{1}$. Exponential and logarithmic maps are defined as
${\mathbf{Exp}^{\mathbb{S}^{D}}_{{\bm{p}}}\left({\bm{u}}\right)}={\bm{p}}\cos({\left\|{\bm{u}}\right\|})+\frac{{\bm{u}}}{{\left\|{\bm{u}}\right\|}}\sin({\left\|{\bm{u}}\right\|})\,,$ (16)
${\mathbf{Log}^{\mathbb{S}^{D}}_{{\bm{p}}}\left({\bm{x}}\right)}=\arccos({\bm{p}}^{T}{\bm{x}})\frac{{\bm{x}}-{\bm{p}}^{T}{\bm{x}}{\bm{p}}}{{\left\|{\bm{x}}-{\bm{p}}^{T}{\bm{x}}{\bm{p}}\right\|}}\,.$ (17)
For the unit sphere $\mathbb{S}^{D}$, the tangent space is Euclidean at any base point, hence the geodesic distance simplifies to
$\text{dist}_{\mathbb{S}^{D}}({\bm{x}},{\bm{p}})={\left\|{\mathbf{Log}^{\mathbb{S}^{D}}_{{\bm{p}}}\left({\bm{x}}\right)}\right\|}=\arccos({\bm{p}}^{T}{\bm{x}})\,.$ (18)
@@
We evaluate RTW in two settings, with ${N}=4$ and ${N}=30$ signals. For transparent results, we measure the computation time on a high-grade GPU and a standard CPU. Furthermore, we evaluate three well-established metrics
@@
where (18) is used within each ${\mathcal{D}_{dtw}}$ distance. The Restoration Accuracy describes how effectively random warping functions are reversed to restore the original signal ${\mathbf{x}_{\text{original}}}$. The Barycenter Loss measures how accurately the mean ${\mathbf{{\hat{y}}}}$ of warped signals ${\mathbf{{\hat{x}}}_{n}}$ represents the data ${\mathbf{x}_{n}}$. The Alignment Quality indicates closeness of aligned signals ${\mathbf{{\hat{x}}}_{n}}$ to their mean ${\mathbf{{\hat{y}}}}$. The average for each metric is computed from 100 experiments.
@@
Metric MMDDTWfootnotemark: TTWfootnotemark: NTWfootnotemark: RTW
Restoration Accuracy 0.0601 0.1712 0.0200 0.0197
Barycenter Loss 0.0552 0.1801 0.0229 0.0206
Alignment Quality 0.0403 0.0878 0.0243 0.0215
@@
Metric p-DTWfootnotemark: TTWfootnotemark: NTWfootnotemark: RTW
Restoration Accuracy 1.0806 0.6458 0.0254 0.0293
Barycenter Loss 2.1259 0.7049 0.2954 0.0689
Alignment Quality 1.2591 0.4332 0.3043 0.0749
CPUfootnotemark: Runtime in [s] 19.5264 18.8754 226.9687 103.1399
GPUfootnotemark: Runtime in [s] 10.6953 9.4830 10.9298 10.2900
@@ V-B1 Evaluation for ${N}=4$
The literature does not offer time warping schemes for ${N}>2$ signals with data on Riemannian manifolds. Hence, we modify MMDDTW [15] to use the Riemannian distance (18), as well as TTW and NTW to adapt (10)-(14). As Table III shows, RTW achieves the best performance across all metrics. While TTW is the fastest, the computational overhead of RTW is well-justified by its results.
@@ V-B2 Evaluation for ${N}=30$
MMDDTW is not capable of handling such a large number of signals in a reasonable time. Instead, we utilize a pairwise approach by iteratively aligning two signals, denoted as p-DTW. The result is shown in Figure 1 and Table IV. RTW performs best across all metrics and the difference to NTW becomes more pronounced.
@@ V-C Multi-dimensional Riemannian manifolds
Next, we extend our experiments to higher-dimensional data by applying RTW to the Riemannian manifolds $\mathbb{S}^{3}$ and $\mathbb{S}^{D}_{++}$.
@@ V-C1 Evaluation in $\mathbb{S}^{3}$
To further validate our results from section V-B, we compare RTW with QDTW on signals with unit quaternions, i.e., $\mathbb{S}^{3}$. Similar to [12], we artificially generate ${N}=4$ valid signals of length ${T}=200$ from the RCFS [27] dataset. This involves projecting handwritten motion data from $\mathbb{R}^{2}$ onto the unit sphere of $\mathbb{S}^{3}$.
In Table V we observe that the mean signal obtained from RTW is an improved representation of the original data compared to QDTW. For the alignment quality, both methods yield similar results, however, RTW is significantly faster.
footnotetext: Modified version with ${N}$-dimensional cost matrix as in MMDDTW [15].
@@
Barycenter Loss 0.0359 0.0339
Alignment Quality 0.0338 0.0338
V-C2 Evaluation in $\mathbb{S}^{D}_{++}$
We temporally align signals from the Riemannian manifold of symmetric positive definite (SPD) $D\times D$ matrices. The exponential and logarithmic maps are
${\mathbf{Exp}^{\mathbb{S}^{D}_{++}}_{{\bm{p}}}\left({\bm{u}}\right)}=\sqrt{{\bm{p}}}\exp{\left(\sqrt{{\bm{p}}}^{-1}{\bm{u}}\sqrt{{\bm{p}}}^{-1}\right)}\sqrt{{\bm{p}}},$ (20)
${\mathbf{Log}^{\mathbb{S}^{D}_{++}}_{{\bm{p}}}\left({\bm{x}}\right)}=\sqrt{{\bm{p}}}\log{\left(\sqrt{{\bm{p}}}^{-1}{\bm{x}}\sqrt{{\bm{p}}}^{-1}\right)}\sqrt{{\bm{p}}},$ (21)
where $\sqrt{{\bm{p}}}$ and $\sqrt{{\bm{p}}}^{-1}$ denote the matrix square root and its inverse, respectively. Note that the Riemannian metric on the SPD manifold is Euclidean only at the identity matrix. Hence, for distance computation under the affine-invariant-Riemannian metric, we use the canonical form where the Frobenius norm ${\left\|\cdot\right\|}_{F}$ yields the geodesic distance
$\text{dist}_{\mathbb{S}^{D}_{++}}({\bm{x}},{\bm{p}})={\left\|\log{\left(\sqrt{{\bm{p}}}^{-1}{\bm{x}}\sqrt{{\bm{p}}}^{-1}\right)}\right\|}_{F}\,.$ (22)
We compare RTW with two modified versions of MMDDTW. The first version, denoted as $\text{MMDDTW}_{{\mathcal{M}}}$, computes the distance between two SPD matrices according to the geodesic distance defined in (22). The second version, denoted as $\text{MMDDTW}_{\operatorname{C}}$, utilizes the Cholesky decomposition $\operatorname{chol}\left(\cdot\right)$ to linearize the manifold, and computes a distance through
@@ V-C2 Evaluation in $\mathbb{S}^{D}_{++}$
The data for this experiment was generated from the RCFS [27] dataset, i.e., a planar robot is controlled to follow the handwritten character G. We create a dataset for $\mathbb{S}^{2}_{++}$ by computing the manipulability as in [28], as shown in Figure 2, as well as $\mathbb{S}^{8}_{++}$ and $\mathbb{S}^{32}_{++}$ by computing the mass matrix for a robot with 8 and 32 links, respectively, resulting in a dataset of ${N}=3$ signals with a length of ${T}=200$ each.
We evaluate the Barycenter Loss (19b) and the Alignment Quality (19c) including the geodesic distance given in (22). Table VI presents the results, supporting our previous findings: RTW outperforms MMDDTW in terms of Barycenter Loss and alignment quality with remarkable speed.
@@
Metric D $\text{MMDDTW}_{{\mathcal{M}}}$ $\text{MMDDTW}_{\operatorname{C}}$ RTW
Barycenter Loss 2 42.1131 43.8988 33.5702
Alignment Quality 2 42.3271 44.6875 23.0621
Barycenter Loss 8 20.8496 21.3290 18.8587
Alignment Quality 8 20.9473 21.4635 16.6931
Barycenter Loss 32 22.2595 23.3155 20.7979
Alignment Quality 32 22.6282 23.5338 18.7008
V-D Real-Robot Teaching with Time Warping in $\mathbb{R}^{3}\times\mathbb{S}^{3}$
Finally, we apply time warping in real-robot teaching scenarios, utilizing a KUKA LBR iiwa manipulator. We demonstrated ${N}=4$ signals through hand-guiding control [29], and recorded the end-effector motion at $30$ Hz, resulting in varying signal lengths ${T_{n}}\in\{2048,\ldots,3667\}$. Subsequently, the mean signal ${\mathbf{{\hat{y}}}}$ obtained from temporal alignment is executed employing a state-of-the-art QP-controller [30], taking hardware limits and manipulator dynamics into account.
For the trajectory teaching, we denote an end-effector pose as ${\bm{x}}=\left[\begin{smallmatrix}{\bm{x}_{\text{pos}}}\\ {\bm{x}_{\text{quat}}}\end{smallmatrix}\right]\in{\left\{\mathbb{R}^{3}\times\mathbb{S}^{3}\right\}}$, consisting of position ${\bm{x}_{\text{pos}}}\in\mathbb{R}^{3}$ and orientation ${\bm{x}_{\text{quat}}}\in\mathbb{S}^{3}$. Similar to [23], we concatenate the logarithmic and exponential map into single operations
${\mathbf{Log}^{\mathbb{R}^{3}\times\mathbb{S}^{3}}_{{\bm{p}}}\left({\bm{x}}\right)}=\left[\begin{matrix}{\bm{x}_{\text{pos}}}\\ {\mathbf{Log}^{\mathbb{S}^{3}}_{{\bm{p}}}\left({\bm{x}_{\text{quat}}}\right)}\end{matrix}\right]=\left[\begin{matrix}{\bm{u}_{\text{pos}}}\\ {\bm{u}_{\text{quat}}}\end{matrix}\right]={\bm{u}}\,,$ (24)
${\mathbf{Exp}^{\mathbb{R}^{3}\times\mathbb{S}^{3}}_{{\bm{p}}}\left({\bm{u}}\right)}=\left[\begin{matrix}{\bm{u}_{\text{pos}}}\\ {\mathbf{Exp}^{\mathbb{S}^{3}}_{{\bm{p}}}\left({\bm{u}_{\text{quat}}}\right)}\end{matrix}\right]=\left[\begin{matrix}{\bm{x}_{\text{pos}}}\\ {\bm{x}_{\text{quat}}}\end{matrix}\right]={\bm{x}}\,.$ (25)
In this experiment, we compare RTW with naive interpolation of the recorded data to the same length, as well as with modified TTWfootnotemark: and NTWfootnotemark: utilizing (24) and (25). Evaluation is done on two tasks, (i) following a constrained path, and (ii) repetitively drawing three circles, both shown in Figure 3.
V-D1 Following a constrained path
We instruct the robot to follow a constrained path with its end-effector, requiring precise control of position and orientation. Due to unintended time shifts within the recorded demonstration data, naive interpolation does not suffice and the resulting mean signal deviates from the desired path, resulting in a collision with the environment. TTWfootnotemark: , NTWfootnotemark: and RTW succeed in following the constrained path without collision, however, the warping functions generated from TTWfootnotemark:  (15) highly compress the motion in certain sections while stretching others, resulting in a mean signal that becomes difficult to execute by the robot, as high velocities are required at the compressed sections.
V-D2 Repetitively drawing three circles
We demonstrate end-effector trajectories consisting of three circles on a whiteboard, which allows us to investigate the time warping performance for repetitive segments within provided signals. Again, naive interpolation is not sufficient. Even though both, TTWfootnotemark: and NTWfootnotemark: , perform better, they do not align all parts of the trajectory properly. By only considering the distance between single points, TTWfootnotemark: and NTWfootnotemark: fail to match repetitive patterns correctly. This results in severe deviations from the desired motion at the start and end of the trajectory for both TTWfootnotemark: and NTWfootnotemark: . Furthermore, TTWfootnotemark: again suffers from fast velocity changes. Only RTW successfully aligns all parts of the demonstration data. Repetitive characteristics of the signal are correctly identified due to the Gaussian-weighted window loss (8) that considers entire segments. When executing the mean signal ${\mathbf{{\hat{y}}}}$ obtained from the proposed RTW approach, the robot’s end-effector successfully draws the three consecutive circles on the whiteboard, as shown in Figure 3.
TABLE VII: Avgerage Reproduction Error (Constrained Path Task)
[…]